a train moving with constant acceleration travels 24ft. during the 10th sec of its motion and 18 ft during the 12th sec of its motion. find its initial velocity and its constant acceleration
1
Expert's answer
2020-02-06T08:23:25-0500

"v=v_0+at"


"24=v_0+10a"

"18=v_0+12a"


We have


"6=-2a \to a=-3 ft/s^2"


"v_0=24-10a=24-10 \cdot (-3)=54 ft/s"


Answer. "a=-3ft\/s^2, v_0=54 ft\/s."